The equivalent resistance of two resistors R1 and R2 is given by π‘…π‘’π‘ž = 𝑅1 + 𝑅2 and π‘…π‘’π‘ž = 𝑅1𝑅2 𝑅1+𝑅2 when they are connected in series and parallel respectively. If 𝑅1 = 22.0 Ξ© Β± 5% π‘Žπ‘›π‘‘ 𝑅2 = 47.0 Β± 5% , find the maximum error in the calculation of their equivalent resistance when they are connected in series and parallel.


1
Expert's answer
2021-11-16T13:46:14-0500

"R'=R_1+R_2" Series combination

Parallel combination

"\frac{1}{R_{eq}}=\frac{1}{R_1}+\frac{1}{R_2}"

Case -1

Error in series combination

"∆R=\pm(∆R_1+∆R_2)"

"∆R=\pm(0.05+0.05)=\pm0.10\Omega"

Now

"R_{eq}=R'\pm∆R"

"R'=22+47=69\Omega"

"R_{eq}=(69\pm0.10)\Omega"

Parallel combination

"\frac{∆R}{R_{eq}}=\frac{∆R_1}{R_1}+\frac{∆R_2}{R_2}"

"\frac{∆R}{R_{eq}}=\frac{0.05}{22}+\frac{0.05}{47}"


"\frac{∆R}{R_{eq}}=(0.002272+0.00106)=0.00333\Omega"

"R_{eq}=\frac{R_1R_2}{R_1+R_2}"

"R_{eq}=\frac{22\times47}{22+47}=14.98\Omega"

"∆R=0.0499\Omega"

"R_{eq}=(14.98\pm0.0499)\Omega"